A driver applies the brakes and slows from 35 m/s to 5 m/s. This takes place over 250 m. How long does it take.
1
Expert's answer
2019-08-28T09:10:18-0400

The distance


"d=\frac{v_i+v_f}{2}t"

The time of braking


"t=\frac{2d}{v_i+v_f}=\frac{2\times 250\:\rm{m}}{(35+5)\:\rm{m/s}}=12.5\:\rm{s}"